Erastic Corporation has $30,000 in cash, $16,000 in marketable securities, $58,000 in account receivable, $72,000 in inventories, and $58,000 in current liabilities. The corporation’s current assets consist of cash, marketable securities, accounts receivable, and inventory. The corporation’s acid-test ratio is closest to:

  • 1.00

  • 1.52

  • 1.79

  • 3.03
Answer #1

Current assets=Cash+Marketable securities+AR+inventory

=30,000+16,000+58,000+72,000=$176000

acid-test ratio=(Current assets-inventory)/Current liabilities

=(176000-72,000)/58,000

which is equal to

=1.79(Approx).